OpenStudy (anonymous):

Estimate seventy-nine point 5 minus 9 point 4 5. Explain your method.

OpenStudy (anonymous):

79.5-9.45=79.50-9.45=70.05

OpenStudy (anonymous):

Thanks! This really helped :3

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

it says estimate. An estimation is not the exact answer. 79.5 - 9.45 (round to the nearest whole number) 80 - 9 = 71 (this is an estimate)

OpenStudy (anonymous):

So dis is where it all started fo you Cat gurl ԅ(◉◞౪◟◉‵ԅ)

OpenStudy (anonymous):

i do not know this .thanks for correcting me.